# From Ambiguity to Verdict: A Semiotic-Grounded Multi-Perspective Agent for LLM Logical Reasoning

### v1

@@ 3.3 Logical Reasoning Stage
Translator.
The translator converts natural-language premises into FOL. Instead of relying on open-ended prompting, this step is guided by a set of general mapping conventions that define how linguistic structures are aligned with logical forms. In particular:
@@
Planner.
For a selected proposition from the semiotic square (e.g., $S_{1}$), the planner constructs a reasoning blueprint. It specifies the evaluation goal, selects relevant premises, and determines which reasoning rules (e.g., Modus Ponens, Modus Tollens, Conjunction, Generalization) could be applied. The planner may also outline potential counterexample checks and identify implicit contextual relations that, while not explicitly stated in the premises, are salient within the discourse background. Its output is a structured reasoning trajectory, but without issuing a verdict.
Solver.
The solver operationalizes the planner’s blueprint: it applies the designated reasoning rules to the given premises, performs deductions step by step, and generates intermediate conclusions. During this process, it verifies logical consistency and checks for contradictions or counterexamples. The solver outputs both a transparent reasoning trace and the final classification of the proposition as True, False, or Uncertain.
@@ 3.4 Reflective Verification Stage
Direct Resolution.
When $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} produce complementary verdicts, such as ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{True}$ and ${\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}=\text{False}$, the stage directly adopts the answer of $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} as final. This scenario reflects a decisive and non-contradictory judgment grounded in the strict contradiction relationship between the proposition and its contradictory. The decision rule for this resolution strategy is defined as follows:
@@
Quick Reflection.
When either $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} or its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is labeled as Uncertain, the stage triggers quick reflection by forwarding the two verdicts and their reasoning traces into a large language model. The model analyzes the internal consistency of the deduction process and returns a refined judgment based on four reflection types:
@@
Deep Reflection.
When both $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} yield the same verdict (e.g., both True or both False), this creates a contradiction under standard logical assumptions. The stage enters Deep Reflection mode, leveraging the structured semantic relations provided by the semiotic square, in particular the implications ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\Rightarrow{\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}$ and ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$, to adjudicate which prediction is more likely to be valid.
@@ 4 RepublicQA Benchmark
Current benchmarks remain largely template-based and semantically simple, offering limited coverage of abstraction and ambiguity. To fill this gap, we construct RepublicQA, a benchmark designed to capture philosophical complexity and multi-perspective reasoning beyond existing datasets.
Benchmark Construction. RepublicQA is built from Plato’s Republic, a foundational text rich in discussions of justice, morality, governance, and knowledge. Its dialogical structure naturally yields abstract propositions and counterarguments, well suited for reasoning tasks involving ambiguity and opposing stances. We extracted propositions and contextual premises manually, with double annotation by two graduate students to ensure logical and semantic consistency.
Complexity Comparison. Unlike benchmarks such as FOLIO, ProofWriter, and ProverQA that remain largely template-based and semantically simple, RepublicQA introduces abstract propositions with deeper logical dependencies, balanced True/False/Uncertain distributions, and contexts requiring the integration of multiple philosophical concepts. We measure complexity across five dimensions: syntactic difficulty (FKGL), vocabulary richness (TTR), lexical variety (MTLD), phrasal diversity (UBR), and vocabulary size. Figure 4 shows that RepublicQA leads on four dimensions, requiring college-level comprehension and exhibiting greater lexical and phrasal diversity, while ranking second in vocabulary size. These properties underscore its abstract nature and resistance to template fitting, making it a valuable benchmark for reasoning under ambiguity and abstraction. More details about the RepublicQA benchmark can be found in Appendix D.
@@ 5.1 Settings
Benchmarks.
We evaluate on four established logical reasoning benchmarks: ProntoQA (Saparov & He, 2022) (5-hop subset), ProofWriter (Tafjord et al., 2020) (depth-5, OWA setting), FOLIO (Han et al., 2022) (full expert-curated split), and ProverQA (Qi et al., 2025) (hard split with 500 examples, 6–9 reasoning steps). Detailed benchmark descriptions are provided in Appendix C.1.
Baselines.
We compare LogicAgent with five representative baselines: Naive Prompting, Chain-of-Thought (CoT) (Wei et al., 2022), Logic-LM (Pan et al., 2023), SymbCoT (Xu et al., 2024b), and Aristotle (Xu et al., 2024a). Detailed baseline introductions are provided in Appendix C.2.
Model.
For RepublicQA, we evaluate with both the locally deployed qwen2.5:32b Yang et al. (2025) and GPT-4o Hurst et al. (2024), ensuring robustness across open and closed source LLMs. For other benchmarks (ProntoQA, ProofWriter, FOLIO, ProverQA), we adopt qwen2.5:32b as the base model. In all experiments, the decoding temperature is fixed at 0.
Symbolic Toolkit.
To verify the syntactic validity of logical forms, we employ the nltk Bird (2006) library for CFG-based structural checking during the FOL parsing stage.
@@
Method RepublicQA Other Benchmarks
Qwen2.5-32B $\uparrow$ GPT-4o $\uparrow$ Avg $\uparrow$ Pronto $\uparrow$ ProofWriter $\uparrow$ FOLIO $\uparrow$ ProverQA $\uparrow$ Avg $\uparrow$
Naive 68.50 74.00 71.25 82.00 59.17 60.29 39.60 60.27
@@
Logic-LM 70.00 73.50 71.75 91.89 63.82 71.93 62.40 72.51
SymCoT 76.00 80.50 78.25 95.20 64.67 70.59 57.20 71.92
Aristotle 74.50 82.50 78.50 94.80 63.23 68.68 56.20 70.73
@@ 5.3 Ablation Study
Q2: Impact of FOL and Natural Language Inputs.
To investigate the respective contributions of FOL and natural language inputs, we conduct an ablation study by removing either the FOL representations or the natural language statements from the model input. The results are visualized in Figure 5(a).
Across four reasoning benchmarks, removing FOL input causes substantial performance degradation. On ProofWriter and FOLIO, accuracy drops from 71.95 to 33.33 and from 79.90 to 34.31, showing that classical FOL representations are critical for symbolic reasoning. Even on ProntoQA, which mainly tests basic deduction, performance falls sharply from 97.80 to 49.00. By contrast, ProverQA shows only a modest decline, likely because its hardest split already involves long reasoning chains, where performance bottlenecks stem more from semantic complexity than from explicit FOL representation alone.
In contrast, RepublicQA shows a different pattern. When the natural language input is removed and only FOL is retained, the accuracy declines substantially from 82.50 to 55.50. Removing FOL leads to a smaller drop to 71.00. This suggests that RepublicQA relies more heavily on the rich semantic information embedded in natural language. As a multi-semantic benchmark grounded in philosophical discourse, it includes contextual nuances such as implicit assumptions, shifting definitions, and pragmatic dependencies, which are not fully captured by symbolic logic alone.
Nevertheless, the best performance on RepublicQA is still achieved when both modalities are used (82.50), confirming that FOL contributes essential symbolic constraints even in semantically complex scenarios. Taken together, these results demonstrate that natural language enhances interpretability and contextual understanding, while FOL ensures inferential precision. Their integration enables more accurate and robust reasoning across a wide range of tasks.
@@ Appendix A Related Work
LLM-Based Logical Reasoning. CoT prompting Wei et al. (2022); Zhang et al. (2025a); Mirzadeh et al. (2024) improves LLM reasoning Gu & Dao (2023); Bai et al. (2023); Huang & Chang (2022) by generating intermediate steps in natural language. Variants such as self-consistency Wang et al. (2022) and symbolic CoT Xu et al. (2024b) improve accuracy and interpretability, with SymbCoT showing that symbolic forms enhance logical reasoning. Aristotle Xu et al. (2024a) further introduces a logic-guided framework using decomposition, search, and resolution, achieving strong results on complex tasks. However, most CoT-based He et al. (2025) and symbolic methods follow linear reasoning paths Zheng et al. (2024); Zhao et al. (2023a) and struggle to capture semantic depth, including contraries and contradictions. Other approaches translate LLM outputs into external logic engines Pan et al. (2023); Ryu et al. (2024), but suffer from brittleness and lack of feedback. In contrast, we propose a fully internal symbolic framework in which the LLM constructs, manipulates, and verifies logical structures. Guided by Greimas’ semiotic square Greimas (1988), our method enables reasoning over semantically diverse, multi-perspective statements.
@@
LLM-Powered Agents. Advances in LLMsSong et al. (2022; 2023); Li et al. (2024); Song et al. (2024); Hu et al. (2025); Song et al. (2025); Ye et al. (2025) have led to agent frameworks Zhang et al. (2025b); Durante et al. (2024) capable of planning, memory, and multi-step reasoning. Systems such as Generative Agents Park et al. (2023), AutoAgents Chen et al. (2023), and MetaGPT Hong et al. (2023) simulate interactive behavior or coordinate task execution, while others like Code-as-Policies Liang et al. (2023), Gorilla Patil et al. (2024), and TaskMatrix Liang et al. (2024) integrate APIs or GUI actions for real-world applications. Building on the agent paradigm, our approach leverages structured and automated reasoning with symbolic representation and reflective verification to tackle abstract and semantically diverse reasoning tasks. By centering the reasoning process on symbolic representation and multi-perspective analysis, we aim to extend the capabilities of LLMs without additional fine-tuning.
@@
Benchmarks for Logical Reasoning. Existing logical reasoning benchmarks Patel et al. (2024) primarily focus on formal validity under controlled conditions. PrOntoQA Saparov & He (2022) is a synthetic benchmark targeting relational reasoning, often centered on transitivity and set membership in fully controlled symbolic settings. ProofWriter Tafjord et al. (2020) features synthetic natural language problems built around symbolic rulebases, where contexts typically involve simplified everyday entities such as animals, colors, and objects, and reasoning proceeds through connectives such as and, or, and not. FOLIO Han et al. (2022) introduces natural language problems annotated with FOL formulas, covering commonsense-style scenarios such as daily events and generic facts. ProverQA Qi et al. (2025) is built with the ProverGen framework, combining LLM generation with automated theorem proving to yield high-quality FOL reasoning tasks. It contains 1,500 dev examples across easy, medium, and hard splits (500 each), with the hard split emphasizing longer reasoning chains and formally verified conclusions. Despite their formal rigor, these benchmarks largely consist of concrete, unambiguous propositions with fixed semantic interpretations. They lack abstract concepts, contextual variability, and structured semantic contrasts such as contraries and contradictions, offering limited support for evaluating multi-perspective, contrastive reasoning over semantically rich and philosophically grounded domains.

### v5

@@ 3.3 Logical Reasoning Stage
Translator. The translator converts natural-language premises into FOL. Instead of relying on open-ended prompting, this step is guided by a set of general mapping conventions that define how linguistic structures are aligned with logical forms. In particular:
@@
Planner. For a selected proposition from the semiotic square (e.g., $S_{1}$), the planner constructs a reasoning blueprint. It specifies the evaluation goal, selects relevant premises, and determines which reasoning rules (e.g., Modus Ponens, Modus Tollens, Conjunction, Generalization) could be applied. The planner may also outline potential counterexample checks and identify implicit contextual relations that, while not explicitly stated in the premises, are salient within the discourse background. Its output is a structured reasoning trajectory, but without issuing a verdict.
Solver. The solver operationalizes the planner’s blueprint: it applies the designated reasoning rules to the given premises, performs deductions step by step, and generates intermediate conclusions. During this process, it verifies logical consistency and checks for contradictions or counterexamples. The solver outputs both a transparent reasoning trace and the final classification of the proposition as True, False, or Uncertain.
@@ 3.4 Reflective Verification Stage
Direct Resolution. When $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} produce complementary verdicts, such as ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{True}$ and ${\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}=\text{False}$, the stage directly adopts the answer of $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} as final. This scenario reflects a decisive and non-contradictory judgment grounded in the strict contradiction relationship between the proposition and its contradictory. The decision rule for this resolution strategy is defined as follows:
@@
Quick Reflection. When either $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} or its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is labeled as Uncertain, the stage triggers quick reflection by forwarding the two verdicts and their reasoning traces into a large language model. The model analyzes the internal consistency of the deduction process and returns a refined judgment based on four reflection types:
@@
Deep Reflection. When both $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} yield the same verdict (e.g., both True or both False), this creates a contradiction under standard logical assumptions. The stage enters Deep Reflection mode, leveraging the structured semantic relations provided by the semiotic square, in particular the implications ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\Rightarrow{\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}$ and ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$, to adjudicate which prediction is more likely to be valid.
@@ 4 RepublicQA Benchmark
Current benchmarks primarily focus on logical complexity while largely overlooking semantic complexity, resulting in limited coverage of abstraction, contextual ambiguity, and nuanced meaning. To address this gap, we construct RepublicQA, a benchmark designed to jointly capture logical depth and semantic breadth reasoning.
Benchmark Construction. RepublicQA draws from classical philosophical and ethical traditions that explore justice, morality, agency, and knowledge. These sources, characterized by dialogical inquiry and abstract argumentation, provide naturally ambiguous propositions and opposing stances suitable for evaluating advanced logical and semantic reasoning. We extracted propositions and contextual premises manually, with double annotation by two graduate students to ensure logical and semantic consistency. A detailed description of RepublicQA and other benchmarks can be found in Appendix C.1.
Complexity Comparison. RepublicQA introduces abstract propositions with deeper logical dependencies, balanced True/False/Uncertain distributions, and contexts requiring the integration of multiple philosophical concepts. To characterize its complexity, we organize our measurements into three categories: a primary indicator reflecting conceptual difficulty (FKGL), secondary indicators capturing lexical and phrasal variation (TTR, MTLD, UBR), and a supporting indicator concerning the structure of contrary construction. Figure 4 shows that RepublicQA achieves the strongest performance across all five complexity indicators, and its contrary construction patterns far exceed those of existing benchmarks, highlighting its emphasis on abstraction, semantic depth, and non-template reasoning. These properties underscore its suitability for evaluating reasoning under ambiguity and high-level conceptual interactions. Details of these metrics are provided in Appendix D.4, with additional information about RepublicQA in Appendix D.
@@ 5.1 Settings
Benchmarks. We evaluate on four established logical reasoning benchmarks: ProntoQA (Saparov and He, 2022) (5-hop subset), ProofWriter (Tafjord et al., 2020) (depth-5, OWA setting), FOLIO (Han et al., 2022) (full expert-curated split), and ProverQA (Qi et al., 2025) (hard split with 500 examples, 6–9 reasoning steps). Detailed benchmark descriptions are provided in Appendix C.1.
Baselines. We compare LogicAgent with five representative baselines: Naive Prompting, Chain-of-Thought (CoT) (Wei et al., 2022), Cumulative reasoning (CR) (Zhang et al., 2023), Tree-of-Thought (ToT) (Yao et al., 2023), Logic-LM (Pan et al., 2023), SymbCoT (Xu et al., 2024b), and Aristotle (Xu et al., 2024a). Detailed baseline introductions are provided in Appendix C.2.
Model. For RepublicQA, we evaluate with both the locally deployed qwen2.5:32b (Yang et al., 2025) and GPT-4o (Hurst et al., 2024), ensuring robustness across open and closed source LLMs. For other benchmarks (ProntoQA, ProofWriter, FOLIO, ProverQA), we adopt qwen2.5:32b as the base model. In all experiments, the decoding temperature is fixed at 0.
Symbolic Toolkit. To verify the syntactic validity of logical forms, we employ the nltk (Bird, 2006) library for CFG-based structural checking during the FOL parsing stage.
@@
Type Method RepublicQA Other Benchmarks
Qwen2.5 $\uparrow$ GPT-4o $\uparrow$ Avg $\uparrow$ Pronto $\uparrow$ ProofWriter $\uparrow$ FOLIO $\uparrow$ ProverQA $\uparrow$ Avg $\uparrow$
LR Naive 68.50 74.00 71.25 82.00 59.17 60.29 39.60 60.27
@@
AR CR 57.00 71.00 64.00 80.20 58.33 71.57 51.80 65.48
ToT 56.00 69.50 62.75 82.50 46.67 72.06 53.40 63.66
SR Logic-LM 70.00 73.50 71.75 91.89 63.82 71.93 62.40 72.51
SymbCoT 76.00 80.50 78.25 95.20 64.67 70.59 57.20 71.92
AR+SR Aristotle 74.50 82.50 78.50 94.80 63.23 68.68 56.20 70.73
@@ 5.3 Ablation Study
  4. 4.
    How do semantic and logical complexity interact? We analyze model performance under varying semantic difficulty and hop depth to make this interplay explicit.
@@
Q2: Impact of FOL and Natural Language Inputs. To assess the individual roles of FOL and natural language inputs, we run ablations that remove either the FOL representations or the natural language statements, as shown in Figure 5(a).
(1) Removing FOL severely impairs symbolic reasoning. Across all four benchmarks, removing FOL causes large accuracy drops, whereas removing natural language leads to only minor degradation. ProofWriter ($71.95\rightarrow 33.33$), FOLIO ($79.90\rightarrow 34.31$), and ProntoQA ($97.80\rightarrow 49.00$). ProverQA shows a smaller decrease ($68.60\rightarrow 60.20$), likely because its natural language descriptions convey relatively clear semantic–logical structure, making it less dependent on explicit FOL representations.
(2) RepublicQA relies more heavily on natural language semantics. When natural language input is removed and only FOL is provided, accuracy decreases substantially $82.50\rightarrow 55.50$, whereas removing FOL leads to a smaller drop to 71.00. This reflects RepublicQA’s higher semantic complexity, which arises from implicit assumptions, shifting definitions, and pragmatic dependencies embedded in philosophical discourse that are not fully recoverable through symbolic logic alone.
(3) Best performance emerges from integrating both modalities. The highest performance across all five datasets is obtained only when both modalities are used, confirming that FOL provides essential symbolic constraints even in semantically complex settings. Overall, natural language contributes interpretability and contextual grounding, while FOL ensures inferential precision; their integration supports more accurate and robust reasoning across diverse tasks.
@@
Table 4: Semantic Difficulty vs. Logical Depth.
[…]
4 0.5161 6.45 0.6029 1.65 0.5636 8.27 0.4286 12.99
5 0.4474 7.72 0.4634 1.60 0.4231 8.40 0.3125 14.68
6 0.3750 7.35 0.4000 1.38 0.2500 7.40 0.0000 17.70
Q4: Semantic–Logical Interplay.
As shown in Table 4, we analyze semantic variation and accuracy trends across different hop depths.
(1) Accuracy uniformly declines with greater logical depth. Across four datasets, accuracy decreases steadily as hop count increases: FOLIO (0.5161 $\rightarrow$ 0.3750), ProofWriter (0.6029 $\rightarrow$ 0.4000), ProverQA (0.5636 $\rightarrow$ 0.2500), and RepublicQA (0.4286 $\rightarrow$ 0.0000).
(2) Prior benchmarks decouple semantic complexity from logical depth. In FOLIO, ProofWriter, and ProverQA, logical depth increases with hop count, but semantic complexity remains nearly constant. For example, FKGL varies only modestly across hops in FOLIO (6.45–7.72), ProofWriter (1.38–1.65), and ProverQA (7.40–8.40).
(3) RepublicQA uniquely increases both semantic and logical complexity. At the same hop level, RepublicQA consistently exhibits higher FKGL yet lower accuracy than all other datasets. For example, around Hop 6, semantic difficulty steadily increases (FKGL $1.38\rightarrow 7.35\rightarrow 7.40\rightarrow 17.70$) while accuracy correspondingly decreases ($0.400\rightarrow 0.375\rightarrow 0.250\rightarrow 0.000$). RepublicQA further shows a clear within-dataset trend: its FKGL rises from 12.99 to 17.70 as hops increase from 4 to 6, while accuracy drops from 0.4286 to 0.0000. This demonstrates that elevated semantic complexity substantially amplifies the difficulty of deep logical reasoning, confirming that RepublicQA is the only benchmark that probes the joint interplay of semantic and logical complexity.
@@ 6 Conclusion
The Usage of LLM
In accordance with ICLR guidelines, we used large language models solely for writing assistance and language refinement.
@@ Appendix A Related Work
LLM-Based Logical Reasoning. CoT prompting (Wei et al., 2022; Zhang et al., 2025a; Mirzadeh et al., 2024) improves LLM reasoning (Gu and Dao, 2023; Bai et al., 2023; Huang and Chang, 2022; Feng et al., 2025b; a) by generating intermediate steps in natural language. Variants such as self-consistency (Wang et al., 2022) and symbolic CoT (Xu et al., 2024b) improve accuracy and interpretability, with SymbCoT showing that symbolic forms enhance logical reasoning. Aristotle (Xu et al., 2024a) further introduces a logic-guided framework using decomposition, search, and resolution, achieving strong results on complex tasks. However, most CoT-based (He et al., 2025; Zhang et al., 2023) and symbolic methods follow linear reasoning paths (Zheng et al., 2024; Zhao et al., 2023a; Sun et al., 2024) and struggle to capture semantic depth, including contraries and contradictions. Other approaches translate LLM outputs into external logic engines (Pan et al., 2023; Ryu et al., 2024), but suffer from brittleness and lack of feedback. In contrast, we propose a fully internal symbolic framework in which the LLM constructs, manipulates, and verifies logical structures. Guided by Greimas’ semiotic square (Greimas, 1988), our method enables reasoning over semantically diverse, multi-perspective statements.
@@
LLM-Powered Agents. Advances in LLMsSong et al. (2022; 2023); Li et al. (2024); Song et al. (2024); Hu et al. (2025); Song et al. (2025); Ye et al. (2025) have led to agent frameworks (Zhang et al., 2025b; Durante et al., 2024) capable of planning, memory, and multi-step reasoning. Systems such as Generative Agents (Park et al., 2023), AutoAgents (Chen et al., 2023), and MetaGPT (Hong et al., 2023) simulate interactive behavior or coordinate task execution, while others like Code-as-Policies (Liang et al., 2023), Gorilla (Patil et al., 2024), and TaskMatrix (Liang et al., 2024) integrate APIs or GUI actions for real-world applications. Building on the agent paradigm, our approach leverages structured and automated reasoning with symbolic representation and reflective verification to tackle abstract and semantically diverse reasoning tasks. By centering the reasoning process on symbolic representation and multi-perspective analysis, we aim to extend the capabilities of LLMs without additional fine-tuning.
@@
Benchmarks for Logical Reasoning. Existing logical reasoning benchmarks (Patel et al., 2024) primarily evaluate formal validity under controlled conditions. PrOntoQA (Saparov and He, 2022) is a synthetic relational reasoning benchmark centered on transitivity and set membership in symbolic settings. ProofWriter (Tafjord et al., 2020) provides synthetic natural language problems grounded in rule-based microworlds with simplified entities and basic logical connectives. FOLIO (Han et al., 2022) contributes natural language scenarios paired with FOL annotations covering everyday commonsense events. ProverQA (Qi et al., 2025), generated via the ProverGen pipeline, combines LLM generation with theorem proving to construct verified reasoning chains across multiple difficulty splits. Despite their rigor, these benchmarks focus almost exclusively on logical form: their propositions are concrete, unambiguous, and semantically fixed. They omit higher-level semantic complexity, including abstract concepts, contextual variability, and systematically constructed relations such as contraries and contradictions. As a result, they offer limited support for evaluating models’ capacity for multi-perspective and contrastive reasoning in semantically rich settings.